# Randomness, Not Representation: The Unreliability of Evaluating Cultural Alignment in LLMs

### v1

@@ 4.2.1. Explicit Value Assessment: Surveys
To directly assess LLM statements about what values they prefer, we use the Value Survey Module (VSM) and Global Opinion Q&A (GQA).
Value Survey Module (VSM)
The VSM survey (also known as Hofstede’s Cultural Survey) consists of 24 standardized questions measuring Hofstede’s cultural dimensions and aggregated human responses from residents of more than 100 countries. We limited our analysis to 65 countries that had complete survey data across all cultural dimensions. To maintain methodological consistency, we adapted the original VSM questions from second-person to third-person format while preserving their semantic meaning. All questions used a 5-point Likert scale (1: Strongly agree, 2: Agree, 3: Undecided, 4: Disagree, 5: Strongly disagree). We calculated cultural dimension values using the VSM Manual’s standardized equations, setting all constants to zero.
Global Opinion QA (GQA)
The GQA dataset offers an additional method for analyzing global opinion distributions across an expanded question corpus, enabling comparisons between LLM responses and country-specific human response patterns. From GQA, we subsampled 180 Likert-scale questions and responses from 15 countries [^2] Brazil, Britain, France, Germany, India, Indonesia, Japan, Jordan, Lebanon, Mexico, Nigeria, Pakistan, Russia, Turkey, USA, selecting only questions with complete responses across all 15 countries. We manually reordered each question’s options to maintain proper Likert scaling.
@@ 4.2.2. Implicit Value Assessment: Job Hiring Sandbox
Cover Letter Evaluation Task
To assess implicit cultural preferences through decision-making scenarios, we designed a job application evaluation task where cultural values significantly influence the selection process. This approach was inspired by the growing adoption of LLMs in recruitment workflows (Kaashoek et al., 2024). We used the ShashiVish/cover-letter-dataset (Vishwakarma, 2023), which comprises 813 cover letters predominantly from technology sector job applications. Using GPT-4o, we generated culturally distinct versions of each cover letter across Hofstede’s six cultural dimensions, creating two polar variants per Hofstede dimension (e.g., Individualism versus Collectivism). We used a standardized prompt for adaptation:
You are a professional cover letter writer. Rewrite the cover letter with a [dimension] tone. Maintain the same content and length of writing as the original cover letter.
This process generated 12 variations per cover letter (two poles for each of Hofstede’s six dimensions). We randomly selected 100 cover letters per dimension for evaluation, yielding 600 comparison pairs. To validate the quality of the generated variations, the authors manually reviewed a random subset of the generated cover letters to verify they accurately reflected the intended cultural dimensions.
@@ 4.3. Statistical Analysis
We used one of two approaches for effect size calculation depending on the number of groups being compared. For binary comparisons, we used the weighted mean difference (WMD), calculated as $|w_{1}\mu_{1}-w_{2}\mu_{2}|$, where $w_{i}$ represents the proportion of samples in group $i$ and $\mu_{i}$ is the group mean. For comparisons involving more than two groups, we implemented a weighted standard deviation (WSD) of means, computed as $\sqrt{\sum w_{i}(\mu_{i}-\bar{\mu})^{2}}$, where $\bar{\mu}$ is the weighted overall mean.
To obtain $p$ values for experiments, we conducted permutation tests with 10,000 iterations, where the observed effect size was compared against a null distribution generated by randomly shuffling ratings between groups while maintaining group sizes. The $p$ value was calculated as the proportion of permuted effect sizes greater than or equal to the observed effect size, with significance levels denoted as * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$). We use this method for all hypothesis tests unless explicitly stated otherwise.
@@ 5.1. Stability
Here, we test stability under both explicit, direct evaluation (in which we ask LLMs questions about culturally relevant values) and an implicit, indirect evaluation (in which we ask LLMs for their preferences between different versions of a cover letter).
@@ 5.1.1. Explicit Preference Evaluation
To test the stability assumption, we examined how sensitive LLM responses are to superficial changes in survey presentation format. Using the GQA dataset, [^3] For this experiment, we further sub-sampled the dataset to only include Likert scale questions with 4 options. we varied two aspects of question presentation that can commonly occur in human survey design (Bogner and Landrock, 2016): Direction (ascending vs. descending order) and Response type (identifier-only vs. full-text responses). For an example of Direction variations, see Table 2, and for an example of Response type variations, see Table 3. We administered these variants across all models, tracking how response patterns shifted with presentation changes while maintaining semantic content.
LLM Responses on VSM vary greatly under trivial (non-semantic) changes to question direction and response type. Figure 2 (Top) shows the mean change in individual responses (0–1 scale) for both Direction and Response changes, revealing that all models display significant shifts when only the presentation type is varied. The overall effect of the variation from simple non-semantic changes to presentation is often larger than real-world cross-cultural differences between humans. Figure 2 (Bottom) shows the overall aggregated change in the evaluation under these changes, demonstrating that these category shifts frequently exceeded the between-country standard deviation benchmark of 0.114.
Figure 2. LLMs’ expressed preferences vary greatly under trivial (non-semantic) changes to question presentation. (Top) Mean absolute valued shifts in individual responses compared to the default prompt for evaluation questions under direction (blue, ascending vs. descending) and response (orange, identifier vs. full text) changes to presentation types. (Bottom) Absolute valued shifts in overall mean evaluation results. The red dashed line shows the standard deviation for actual human responses (0.114). The overall change in assessed preference often exceeds one human standard deviation. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001. (Top) Chi-square test against the null hypothesis that response categories are independent of format changes. (Bottom) One-sided permutation test (10,000 iterations) against the null hypothesis that shifts in model outputs between format conditions are due to random chance.
@@ 5.1.2. Implicit Preference Evaluation
In contrast to the above where we directly asked LLMs for their preferences about values, here we test LLMs’ stability under evaluation of their implicit biases by having them compare and rate different versions of cover letters (see Section 4.2.2).
@@
Comparative/Absolute Assessment Instability
First, we compared comparative versus absolute (non-comparative) approaches for having LLMs assess cover letters. In the comparative setting, we asked an LLM for a Likert scale preference between two examples, while in the absolute one, we asked them for Likert scale ratings of individual examples. This is motivated in part by systematic differences in human judgments between joint and separate evaluation modes (Hsee, 1996). See Table 4 for the prompts that we used for both the comparative and absolute rating assessments. To enable direct comparison, we normalized both scales to a [-1, 1] range. Comparative ratings were centered and scaled to maintain their inherent bipolar nature, while pairs of absolute ratings were aggregated by taking the normalized differences between the two ratings. This approach preserves both preference direction and strength—negative values indicating low-dimension preference and positive values high-dimension preference—allowing us to systematically examine how rating methods influence cultural dimension assessments.
@@
Comparative versus absolute preference elicitation questions yield different cultural preferences from LLMs. Figure 3 shows systematic differences in score distributions: comparative ratings generally show wider variance and more extreme negative values, particularly for Indulgence/Restraint and Long/Short Term Orientation, while absolute ratings tend to cluster more tightly around neutral values.
Figure 3. LLM cover letter evaluations vary greatly under comparative versus absolute preference elicitation questions. Normalized preferences (-1 to +1) for comparative (blue) and absolute (orange) ratings. Comparative ratings show greater variance, especially for Indulgence/Restraint and Long-term Orientation. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to a one-sided permutation test with 10,000 iterations against the null hypothesis that there is no difference in mean preference scores between comparative and absolute rating conditions.
Scale Instability
Testing different Likert scale configurations helps determine whether LLM judgments maintain proportional relationships across rating resolutions. Social science researchers have long established that the design of Likert scales can influence humans’ ratings (Krosnick and Fabrigar, 1997). Here, we examine whether LLM judgments shift when moving from a 4-point to a 6-point format. For examples of the different rating scales used, see Table 5.
Likert scale size affects LLMs’ implicit cultural preferences. Figure 4 reveals substantial distributional differences in ratings across scales. These variations are quantified in Table 1, where Indulgence/Restraint and Long/Short Term Orientation show the strongest sensitivity to scale changes (effect sizes of 0.088 and 0.067 respectively, $p<0.001$), with responses shifting from concentrated patterns in the 4-point scale to more dispersed or extreme patterns in the 6-point scale. In contrast, Masculinity/Femininity maintains relatively consistent response patterns across different scales.
Figure 4. The Likert scale size affects LLM’s implicit cultural preferences. Response patterns across 4/5/6-point scales. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to a one-sided permutation test with 10,000 iterations against the null hypothesis that there is no difference in the mean normalized ratings between different Likert scale sizes.
Context Instability
Research has shown that humans systematically attribute different levels of knowledge to individuals based solely on their contextual roles, even when given identical information about them (Ross et al., 1977). This demonstrates how context can shape human evaluation of unchanged content. To examine whether language models exhibit context-dependent evaluation patterns, we assessed the same cover letters from three distinct professional perspectives: Hiring Manager, Job Applicant, and Career Coach (for complete prompts, see Table 6).
Trivial variations in the role that an LLM is prompted to play can affect LLM cultural preferences. Figure 5 reveals systematic variations across trivial variations in the role that an LLM is prompted to play. Long/Short Term Orientation shows the most dramatic context sensitivity (mean 0.367, $p<0.001$), with Job Applicants rating significantly lower than both Hiring Managers and Career Coaches. Uncertainty Avoidance (mean 0.338, $p<0.001$) follows a similar pattern, while Individualism/Collectivism (mean 0.293, $p<0.001$) shows an inverse relationship where Job Applicants and Career Coaches rate consistently higher than Hiring Managers. In contrast, dimensions like Indulgence/Restraint (mean 0.064), Masculinity/Femininity (mean 0.088), and Power Distance (mean 0.106) maintain more consistent ratings across contexts (all $p>0.05$).
Figure 5. Trivial changes in the role that an LLM is prompted to play can influence expressed cultural preferences. Largest variations in Long-term Orientation (3.74 to 2.92) and Uncertainty Avoidance (2.84 to 2.01). Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to a one-sided permutation test with 10,000 iterations against the null hypothesis that there is no difference in the mean ratings between different roles.
Figure 6. Reasoning requirements alter rating distributions. Rating patterns with reasoning (blue) and without reasoning (orange) shown across cultural dimensions. Most dramatic shifts in Indulgence/Restraint (from rating 1 to 4) and Power Distance (from dispersed to concentrated). Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to a one-sided permutation test with 10,000 iterations against the null hypothesis that there is no difference in mean ratings between conditions with and without reasoning prompts.
Figure 7. Extrapolation across cultural dimensions is unreliable in humans and LLMs alike. The validity of extrapolation is highly sensitive to the geometry of individual cultural dimensions (Top) For humans and LLMs, extrapolability (as measured by clustering ARI) increases with the number of observed dimensions. However, it is near a random guess baseline for low numbers of observed dimensions. (Bottom) Different cultural dimensions have very different impacts on the validity of extrapolation between dimensions. For countries (blue), Indulgence/Restraint strengthens groupings while Masculinity/Femininity weakens them. For LLMs (orange), Uncertainty Avoidance Index strengthens while Long/Short Term orientation weakens the clustering. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001. (Top) For each group (Countries/LLM), the null hypothesis was that the clustering similarity between subsets of dimensions and all dimensions could arise from random cluster assignments. (Bottom) For each group (Countries/LLM), the null hypothesis was that there is no difference in mean ARI scores between dimension subsets that include versus exclude each individual dimension.
Reasoning Instability
A request to justify decisions with reasoning can affect human judgment (Wilson and Schooler, 1991), and humans often construct rationales after making decisions (Haidt, 2001). To examine whether requiring explicit reasoning affects evaluation stability, we compared ratings with and without asking LLMs for explanations before their answer (for complete prompts, see Table 7).
@@
Asking for reasoning affects LLMs’ expressed cultural preferences. Table 1 and Figure 6 reveal large shifts between direct rating and reasoning-required ratings. Long/Short Term Orientation exhibits the largest effect (mean 0.770, $p<0.001$), with the histogram showing how the model clusters around rating 4 without reasoning, but disperses more evenly across ratings 1-5 when required to explain its choices. Similarly, Indulgence/Restraint shows substantial sensitivity (mean 0.487, $p<0.001$), with ratings shifting from a strong concentration at rating 1 to a preference for rating 4 when explanations are needed. Masculinity/Femininity (mean 0.283, $p<0.001$) also shows significant redistribution in the histogram, while other dimensions show changes with relatively high $p$ values ($p>0.05$).

### v2

@@ 4.2.1. Explicit Value Assessment: Surveys
To directly assess LLM statements about what values they prefer, we used two established surveys: Value Survey Module (VSM) and Global Opinion Q&A (GQA).
Value Survey Module (VSM): The VSM survey (also known as Hofstede’s Cultural Survey) consists of 24 standardized questions measuring Hofstede’s cultural dimensions. It also contains aggregated human responses from residents of more than 100 countries. We limited our analysis to include only the 65 countries that had complete survey data across all cultural dimensions. To maintain methodological consistency, we adapted the original VSM questions from second-person to third-person format while preserving their semantic meaning. Each question used a consistent 5-point Likert scale (1: Strongly agree, 2: Agree, 3: Undecided, 4: Disagree, 5: Strongly disagree). Cultural dimension scores were calculated using the VSM Manual’s standardized equations with all constants set to zero. This survey was used in our extrapolability experiments (see Section 5.2).
Global Opinion QA (GQA): The GQA dataset offers an additional method for analyzing global opinion distributions across an expanded question corpus. We filtered for questions with Likert-scale response options, selecting 180 questions that had complete response data across 15 countries (Brazil, Britain, France, Germany, India, Indonesia, Japan, Jordan, Lebanon, Mexico, Nigeria, Pakistan, Russia, Turkey, USA). All selected questions used a 4-point Likert scale (1: Strongly agree, 2: Agree, 3: Disagree, 4: Strongly disagree). Since the original dataset contained inconsistent orderings, we standardized all response options to ensure consistent directionality across questions. This standardization enabled valid comparisons between LLM responses and country-specific human opinions. We used this survey in both our explicit stability experiments (see Section 5.1.1) and steerability experiments (see Section 5.3).
@@ 4.2.2. Implicit Value Assessment: Job Hiring Sandbox
To assess implicit cultural preferences, we developed a job application evaluation task inspired by LLMs’ growing use in recruitment (Kaashoek et al., 2024). This approach reveals cultural biases through hiring decisions rather than direct questioning about values, providing ecological validity.
Using the ShashiVish/cover-letter-dataset (Vishwakarma, 2023) of 813 technology sector cover letters, we generated culturally distinct variants along Hofstede’s six dimensions. For each original letter, GPT-4o (using a temperature of 1.0 to ensure sufficient variation) created contrasting pairs representing polar cultural values (e.g., Individualism vs. Collectivism) using the prompt: ”You are a professional cover letter writer. Rewrite the cover letter with a [dimension] tone. Maintain the same content and length of writing as the original cover letter.” We randomly selected 100 cover letters per dimension (600 comparison pairs total) and asked LLMs to express preferences between these variants using different evaluation formats as described in the implicit stability experiment (see Section 5.1.2).
@@ 4.3. Statistical Analysis
We applied the following statistical procedures across all experiments to ensure consistent analysis of cultural alignment.
Normalization: To enable valid comparisons across different Likert scales, we normalized responses to a continuous scale using: $normalized\_rating=(raw\_rating-min\_rating)/(max\_rating-min\_rating)$, resulting in values from 0 to 1 (e.g., a rating of 3 on a 1-5 scale becomes 0.5).
For comparative versus absolute evaluations in the stability experiments, we instead normalized to a [-1, 1] range to preserve directionality. Comparative ratings were centered and scaled to maintain their bipolar nature, while pairs of absolute ratings were aggregated as normalized differences to preserve both preference direction and magnitude.
Effect Size: To quantify the magnitude of differences between experimental conditions independent of sample size, we calculated effect sizes. For binary experimental conditions (comparative/absolute format and reasoning/non-reasoning requirements), we used weighted mean difference (WMD): $|w_{1}\mu_{1}-w_{2}\mu_{2}|$, where $w_{i}$ represents the proportion of samples in condition $i$ and $\mu_{i}$ is the condition mean (e.g., the absolute difference between average normalized ratings in comparative versus absolute conditions). For multi-category conditions (4/5/6-point Likert scales and Hiring Manager/Job Applicant/Career Coach contexts), we used weighted standard deviation (WSD) of means: $\sqrt{\sum w_{i}(\mu_{i}-\bar{\mu})^{2}}$, where $\bar{\mu}$ is the weighted overall mean.
Hypothesis Testing: To determine statistical significance of observed differences, we obtained p-values through permutation tests with 10,000 iterations, comparing observed effect sizes against null distributions generated by randomly shuffling ratings between experimental conditions while maintaining group sizes. The $p$ value represents the proportion of permuted effect sizes greater than or equal to the observed effect size, with significance levels denoted as * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$). We conducted separate tests for each cultural dimension to isolate dimension-specific effects.
@@ 5.1. Stability
Here, we test stability under both explicit, direct evaluation (by asking LLMs questions about culturally relevant values) and an implicit, indirect evaluation (by asking LLMs for preferences between different versions of a cover letter).
@@ 5.1.1. Explicit Preference Evaluation
We examined how sensitive LLM responses are to superficial changes in survey presentation format. Using the GQA dataset, we varied two aspects of question presentation that can commonly occur in human survey design (Bogner and Landrock, 2016): (1) Direction — whether response options appear in ascending order (e.g., 1=”Very important” to 4=”Not at all important”) versus descending order (e.g., 1=”Not at all important” to 4=”Very important”), and (2) Response type — whether LLMs must respond with only numerical identifiers (e.g., ”2”) versus full text options (e.g., ”Rather important”). For complete examples of these prompts, see Table 2 and Table 3.
After varying both Direction and Response type formats, we analyzed the resulting LLM responses in two ways. We measured (1) the normalized category shift size, which reflects the proportion of questions where LLMs changed their response solely due to presentation format changes, and (2) the effect size (using Weighted Mean Difference) of these non-semantic changes, measuring the average magnitude of shifts across all questions. For this analysis, we compared effect sizes against a human baseline, the standard deviation of between-country differences (0.114) in the GQA dataset.
LLM Responses on the GQA survey vary greatly under non-semantic changes to options direction and response type. Figure 2 (Left) shows the normalized category shift size for both Direction and Response Format. All models show significant changes when only the presentation format is varied. Figure 2 (Right) shows the effect size (Weighted Mean Difference) of these presentation variations, demonstrating that changes frequently exceeded the between-country standard deviation benchmark of 0.114. This indicates that superficial format changes often produce larger effects than real-world cross-cultural differences between humans.
Figure 2. LLMs’ expressed preferences vary greatly under non-semantic changes to question presentation. (Left) Normalized Category Shift Size shows proportion of maximum possible shift when changing Direction Format (blue, ascending vs. descending) or Response Format (orange, identifier-only vs. option-text). (Right) Effect Size (Weighted Mean Difference) measures response magnitude changes between format conditions. Red dashed line represents one standard deviation (0.114) of between-country human response variation. The overall change in assessed preference often exceeds one human standard deviation. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001. (Left) Chi-square test against the null hypothesis that response categories are independent of format changes. (Right) One-sided permutation test with 10,000 iterations against the null hypothesis that shifts in model outputs between presentation conditions are due to random chance.
@@ 5.1.2. Implicit Preference Evaluation
In contrast to Section 5.1.1 where we directly asked LLMs for their preferences about values, here we test LLMs’ stability under evaluation of their implicit biases by having them compare and rate different versions of cover letters (see Section 4.2.2). This approach reveals cultural preferences in decision-making behavior rather than through stated values.
We systematically change four aspects of the evaluation methodology: (1) comparative versus absolute formats, (2) reasoning requirements, (3) Likert scale design, and (4) contextual framing. These modifications were chosen because they deliberately mirror methodological manipulations commonly used in human survey research to test preference stability. For each modification, we analyze rating distributions across cultural dimensions and quantify effect sizes using Weighted Mean Difference (WMD) for binary comparisons and Weighted Standard Deviation (WSD) for multi-category comparisons. We use permutation tests to determine whether these changes produce statistically significant differences in expressed preferences.
@@
Comparative/Absolute Assessment Instability: We juxtaposed comparative versus absolute (non-comparative) approaches to cover letter assessment. In the comparative setting, we asked an LLM to rate its preference between two letters on a 5-point Likert scale (e.g., ”Strongly prefer Cover Letter A” to ”Strongly prefer Cover Letter B”), while in the absolute setting, we presented the same pairs but asked for independent ratings of each letter on a similar 5-point scale (e.g., ”Not likely at all” to ”Very likely” to select the candidate). This approach is motivated in part by systematic differences in human judgments between joint and separate evaluation modes (Hsee, 1996). See Table 4 for the prompts used.
@@
Comparative versus absolute preference elicitation questions yield different cultural preferences from LLMs. Table 1 and Figure 3 (Left) shows systematic differences in score distributions: comparative ratings generally show wider variance and more extreme negative values while absolute ratings tend to cluster more tightly around neutral values.
Reasoning Instability: We evaluated cover letter assessments with and without reasoning requirements. In the standard condition, LLMs directly provided ratings without explanation, while in the reasoning condition, they were instructed to provide their rationale before giving a numerical rating. This manipulation isolated the impact of explanation requirements while maintaining identical cover letter content and rating scales. This experiment is motivated by evidence that requiring justification can affect human judgment (Wilson and Schooler, 1991), and that humans often rationalize decisions post-hoc (Haidt, 2001). We sought to determine whether LLMs exhibit similar shifts in judgment when prompted to explain their reasoning before rating (for complete prompts, see Table 7).
@@
Asking for reasoning affects LLMs’ expressed cultural preferences. Table 1 and Figure 3 (Right) reveal large shifts between direct rating and rating with reasoning. These shifts are particularly pronounced in Long/Short Term Orientation and Indulgence/Restraint preferences, with significant redistribution across rating categories when explanations are required.
Figure 3. Binary variations in evaluation design impact LLMs’ expressed cultural preferences. (Left) LLM cover letter evaluations vary under comparative versus absolute preference elicitation. Normalized preferences (-1 to +1) for comparative (blue) and absolute (orange) ratings reveal differences in distributions across cultural dimensions. (Right) Reasoning requirements alter rating distributions. Rating patterns with reasoning (blue) and without reasoning (orange) show varying distributions across the same dimensions. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to one-sided permutation tests with 10,000 iterations against the null hypothesis of no difference in mean ratings between conditions.
Scale Instability: We compared cover letter evaluations across three Likert scale formats: 4-point, 5-point, and 6-point scales. Each maintained the same response pattern, varying only in available options. This experiment is motivated by the long established findings that the design of Likert scales can influence humans’ ratings (Krosnick and Fabrigar, 1997). For examples of the scales used, see Table 5.
Likert scale size affects LLMs’ implicit cultural preferences. Figure 4 (Left) reveals substantial distributional differences in ratings across scales. These variations are quantified in Table 1, with responses generally shifting from more concentrated patterns in the 4-point scale to more dispersed or extreme patterns in the 6-point scale.
Context Instability: We prompted LLMs to evaluate cover letters under three different personas: Hiring Manager, Job Applicant, and Career Coach. These roles were selected because, while distinct, they should not elicit different views on cover letter quality. Only the role description in the prompt was modified, with all other elements remaining identical. This approach tested whether LLMs maintain consistent cultural preferences despite minimal contextual variations, which would suggest robust internal value representations. Our experiment was informed by findings that humans sometimes attribute different levels of knowledge to individuals based solely on their contextual roles, even when given identical information about them (Ross et al., 1977) (for complete prompts, see Table 6).
Trivial variations in the role that an LLM is prompted to play can affect LLM cultural preferences. Table 1 and Figure 4 (Right) reveal systematic variations across different prompted roles, with Long/Short Term Orientation and Uncertainty Avoidance showing the most pronounced context sensitivity, while dimensions like Indulgence/Restraint, Masculinity/Femininity, and Power Distance remained more consistent.
Figure 4. Multi-category variations in evaluation design affect LLMs’ cultural preferences. (Left) The Likert scale size affects LLM’s implicit cultural preferences. Response patterns across 4-point (blue), 5-point (orange), and 6-point (green) scales show differences in preference distributions across cultural dimensions. (Right) Trivial changes in the role that an LLM is prompted to play can influence expressed preferences. Normalized ratings from Hiring Manager (blue), Job Applicant (orange), and Career Coach (green) perspectives show systematic variations. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001 according to one-sided permutation tests with 10,000 iterations against the null hypothesis that there is no difference in mean ratings between conditions.